OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x − 1) = 5x − (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

4 + 2(x - 1) = 5x - (x + 10) 4 + 2x - 2 = 5x - x - 10 2x - 5x + x = - 4 + 2 - 10 -2x = -12 x = -12/-2 x = 6